# Secure Video Quality Assessment Resisting Adversarial Attacks

### v1

@@ Abstract
The exponential surge in video traffic has intensified the imperative for Video Quality Assessment (VQA). Leveraging cutting-edge architectures, current VQA models have achieved human-comparable accuracy. However, recent studies have revealed the vulnerability of existing VQA models against adversarial attacks. To establish a reliable and practical assessment system, a secure VQA model capable of resisting such malicious attacks is urgently demanded. Unfortunately, no attempt has been made to explore this issue. This paper first attempts to investigate general adversarial defense principles, aiming at endowing existing VQA models with security. Specifically, we first introduce random spatial grid sampling on the video frame for intra-frame defense. Then, we design pixel-wise randomization through a guardian map, globally neutralizing adversarial perturbations. Meanwhile, we extract temporal information from the video sequence as compensation for inter-frame defense. Building upon these principles, we present a novel VQA framework from the security-oriented perspective, termed SecureVQA. Extensive experiments indicate that SecureVQA sets a new benchmark in security while achieving competitive VQA performance compared with state-of-the-art models. Ablation studies delve deeper into analyzing the principles of SecureVQA, demonstrating their generalization and contributions to the security of leading VQA models.
@@ I Introduction
The tremendous surge of streaming media platforms like YouTube, Facebook, and TikTok has resulted in an exponential increase in video traffic [1]. This phenomenon has transformed ordinary individuals into both providers and end-users of video services. There is an essential need for Video Quality Assessment (VQA) to automatically evaluate the quality of videos and detect videos that are not up to standard. In general, VQA is categorized into three types based on the availability of the pristine video: Full-Reference VQA (FR-VQA), Reduced-Reference VQA (RR-VQA), and No-Reference VQA (NR-VQA). Due to the characteristics of videos on the Internet, NR-VQA has garnered the widest attention [36, 37, 40]. In recent years, driven by the development of cutting-edge architectures, NR-VQA has experienced rapid advancements and achieved human-comparable performance [41, 42, 18, 43].
@@
With the above considerations, this paper makes the first attempt to investigate some general adversarial defense principles that can be applied to existing VQA models in a plug-and-play manner and design a robust NR-VQA framework from a security standpoint, termed SecureVQA. It is worth noting that SecureVQA does not simply adopt the defense strategies used in classification. Instead, it ingeniously devises solutions tailored to the characteristics of VQA task. In summary, the key idea behind the proposed framework is to weaken the adversarial effect by subtly introducing some randomness and a part of additional information as a guarantee. Specifically, SecureVQA designs two branches and applies three defense strategies to each branch for this purpose, each of which can serve as a general component for existing VQA models. Firstly, we empirically deem randomness as the natural enemy against optimization-based adversarial attacks. Based on it, we introduce random spatial grid sampling on the video frame for intra-frame defense, which can improve the model’s robustness while preserving the detailed information of the video frame. Secondly, a universal strategy termed pixel-wise randomization is proposed. The pixel-wise randomization deliberately introduces random imperceptible modifications through a guardian map to the video before assessment, effectively destroying the carefully crafted perturbations while maximally avoiding the loss of intra-frame information of the video. Thirdly, incorporating additional temporal information is also crucial for model robustness. Given that current attack methods primarily focus on introducing spatial perturbations, the distinctive temporal information of videos serves as an additional safeguard for the model’s security, which also avoids additional modifications to the inter-frame information. The objective of SecureVQA against adversarial attacks is shown in Fig. 1. It is anticipated that our framework can not only have strong anti-interference ability for carefully designed adversarial videos but also obtain superior VQA performance. The major contributions of this work are summarized as follows:
@@ II-A NR-VQA Models
Since [17] first introduced deep learning into NR-VQA, significant progress has been made in improving the capabilities of models from multiple perspectives [34, 35, 41, 42, 18]. [49] considered the influence of motion distortion and processed the video into continuous chunks, enabling the extraction of spatial-temporal features from the video. [23] introduced texture information as a complementary component to content features, aiming to compensate for the human eye’s tendency to overlook distortions that occur in regions with intricate textures. [20] incorporated a blur encoder into the model to detect the blur effect, recognizing the significant influence of video shakiness on quality. [22] discovered that abrupt scene transitions result in additional degradation in quality, and the Human Visual System (HVS) tends to allocate different levels of attention to frames with varying contents. Subsequently, two Transformer-based modules were designed to handle these two characteristics. [24] systematically divided the video into 13 dimensions to perform a detailed analysis of the individual factors influencing video quality. Additionally, a vision-language-based model was designed to jointly predict specific quality factors and overall quality scores. [19] conducted a thorough survey of HVS, and designed five modules to mimic the five key characteristics of HVS. [21] partitioned the VQA model into technical and aesthetic branches, allowing for a comprehensive evaluation of video quality from two distinct perspectives. [11] explored the potential of Large Language Models (LLMs) in VQA tasks for the first time. However, the results indicated that it was extremely difficult for LLMs to output accurate scores directly. Subsequently, [12] transformed score regression into classification problem and used the weighted sum as the output scores, validating the ability of LLMs in VQA.
@@ II-B Adversarial Attacks on QA models
In the past two years, some studies have revealed that similar to classification models, QA models are also vulnerable to adversarial attacks [2, 3, 4, 10, 15, 16]. [2] made the first attempt to attack the NR-IQA models under the white-box setting. The perturbation was randomly initialized from a discrete set. During the optimization process, the steepest ascent method was utilized to calculate the gradient and update the perturbation at each iteration. To ensure the visual quality of the adversarial images, [2] performed human subjective experiments to select the adversarial image that is most similar to the original one. [3] aimed to increase the quality scores of all QA models by training a fixed-size universal perturbation. In their loss function, a normalization factor was introduced to balance the trade-off between the attack effect and the visual quality. [4] conducted the first endeavor to attack the NR-VQA models under the black-box setting, where the attack problem was formulated as misleading the adversarial video’s estimated quality score far away from its ground-truth score towards a specific boundary. Recognizing the time-consuming and labor-intensive human subjective experiments, they employed objective metrics $L_{2}$ and $L_{\infty}$ norm to constrain the magnitude of perturbation automatically. Besides, [9] proposed a black-box adversarial attack against NR-IQA models through a random search method in a greedy way. [13] proposed the correlation-error-based attack framework, which specifically perturbed the score variation on individual images and the correlation within the image set. [7] proposed a query-based black-box attack method against NR-IQA models with adaptive iterative attacks guided by the initial attack direction.
@@ III-A Intra-frame defense
which can effectively eliminate the carefully crafted adversarial perturbation while preserving the visual quality of the defense video with minimal alterations. After applying the spatial sampling and pixel-wise randomization strategies, the input video $\mathbf{V}$ is transformed into $\mathbf{F}^{\text{intra}}$ in the intra-frame branch, and then $\mathbf{F}^{\text{intra}}$ are fed into video Swin-Transformer (Swin-T) to obtain intra-embeddings $\mathbf{E}^{\text{intra}}$ as
@@ III-B Inter-frame defense
where $s:s+d$ denotes selecting $d$ consecutive frames for extracting temporal information, and $s$ is set to a different value from the intra-frame branch. Considering that the details and high-frequency noise of the video frame are preserved through spatial grid sampling in the intra-frame branch, we therefore, employ the resize operation as a supplementary in the inter-frame branch to comprehensively extract video embeddings by
@@
Similarly, the pixel-wise randomization is also adopted through guardian map to introduce randomness and neutralize the meticulously crafted perturbation due to its minimal impact on performance and significant contribution to security by
@@
Temporal information extraction: After applying the resize operation and pixel-wise randomization strategies, the input video $\mathbf{V}$ is transformed into $\mathbf{F}^{\text{inter}}$ in the inter-frame branch, and then $\mathbf{F}^{\text{inter}}$ are fed into SlowFast by
@@
TABLE I: Performance evaluations under white-box setting on VSFA [17], MDTVSFA [40], TiVQA [23], BVQA-2022 [38], and HVS-5M [19]. Here, the performance before attack is marked in gray.
@@
TABLE II: Performance evaluations under black-box setting on VSFA [17], MDTVSFA [40], TiVQA [23], BVQA-2022 [38], and HVS-5M [19]. Here, the performance before attack is marked in gray.
@@ IV-A Experimental setup
Attack strategy: Currently, research on the security of VQA is in its preliminary stage. Most of the related work focus on adversarial attacks on IQA. As far as we know, [4] is the only representative adversarial attack on VQA under the black box setting. Therefore, we select [4] as the white-box and black-box attack benchmark in this paper, which designs an efficient way to address the computation issue within adversarial video. Specifically, [4] employs the Projected Gradient Descent (PGD) [50] algorithm for white-box attack, restricting the pixel-wise $L_{2}$ or $L_{\infty}$ norm of the perturbations to 1/255 or 3/255 in each iteration. In black-box attack, [4] tentatively perturbs a portion of the frame (56$\times$56 pixels) while leaving the remainder untouched at each query. Such perturbations would be keep if the model’s score changes in the expected directions, and abandoned otherwise. After 300 times of queries, the resulting adversarial frame is successfully obtained. Besides, considering the computational complexity, [4] selects 50 videos in the dataset to verify the model security. In this paper, we adopt the same setting as [4] for fair comparison.
Implementation details: The experiments are built on Pytorch 3.8 and run on one RTX 4090 GPU card. In the two branches that handle the intra-frame and inter-frame information, the Swin-Transformer [25] pre-trained on ImageNet [39] and SlowFast [30] pre-trained on Kinetics [31] are utilized as backbone networks to extract embeddings. Among them, the parameters of the Swin-Transformer are trainable, while the parameters of the SlowFast are frozen. In the skip sampling and continuous sampling process, $n$ and $d$ (in Equations 1 and 7) are set to 2 and 32 respectively, and $\mathbf{V}_{t}^{\text{inter}}$ is resized to 224 $\times$ 224 (in Equation 8). In addition, the inter-frame branch splits the video into 16 segments. In the spatial sampling process, $G$ and $S$ (in Equations 2 and 3) are set to 7 and 32, respectively, as done in [18]. The Adam [32] optimizer with the initial learning rate of 0.001 and batch size 12 is used for training.
TABLE III: Evaluations on the large-scale dataset. Here, the top three performers are highlighted in boldface, underlined, and wavy line respectively, in descending order of performance.
@@
TABLE IV: Evaluations on smaller datasets. Here, the top three performers are highlighted in boldface, underlined, and wavy line respectively, in descending order of performance.
@@ IV-B Evaluations on model security
To further validate the security of SecureVQA and its ability to maintain the imperceptibility of the perturbations in the original video, we present some examples of the predicted scores of the original video frames and the corresponding adversarial video frames in Fig. 3. It can be observed that SecureVQA offers formidable security against both white-box and black-box attacks. Specifically, SecureVQA not only prevents high quality videos from being rated as low quality, it also works the other way around. Furthermore, the modifications introduced by SecureVQA are virtually undetectable, preserving the detailed information and the quality of the video content to the fullest extent.
TABLE V: Ablation studies on KoNViD-1k. Here, the performance before attack is marked in gray, and the performance improvement and degradation are colored in red and blue, respectively.
@@
TABLE VI: Ablation studies on LIVE-VQC. Here, the performance before attack is marked in gray, and the performance improvement and degradation are colored in red and blue, respectively.
@@ IV-C Evaluations on quality prediction
In this part, the performance of SecureVQA on clean videos is evaluated on the large-scale LSVQ dataset. Besides, it is further fine-tuned on the small-scale datasets, and the results are given in Tables III and IV, respectively. It can be seen that SecureVQA achieves competitive performance compared with SOTA VQA models. Specifically, SecureVQA has consistently secured a position within the top three across all test categories. Meanwhile, its remarkable performance on the LIVE-VQC dataset which encompasses rich motion information, can be attributed to SecureVQA’s meticulous consideration of inter-frame information and its effective fusion of embeddings from two distinct branches. Considering that SecureVQA takes security as the prerequisite and then ensures its exceptional VQA performance. As a result, certain strategies like the guardian map may have a detrimental effect on performance due to the inherent tradeoff between security and accuracy (as shown in the ablation study). Despite this, SecureVQA exhibits an accuracy difference of merely 1% to 2% compared to the most advanced DOVER [21]. The results from both this part and the previous part confirm that SecureVQA has successfully achieved the expected goal depicted in Fig. 1. It not only demonstrates a formidable resistance against adversarial attacks but also attains superior VQA performance.
@@ IV-D Ablation study
Figure 4: The defense effect of guardian map added in different regions and corresponding predicted scores. Here, the pink area denotes the size of the disturbed area after the attack.
Firstly, the generality of these three strategies to existing models is verified. Each strategy is individually incorporated into the original model in a plug-and-play manner to verify its defense effect and characteristics. Specifically, 1) Inter-frame embeddings can be easily concatenated with the embeddings of existing models. 2) The existing models typically feed pre-processed frames into the networks, and we simply need to substitute them to the frames processed by spatial grid sampling. 3) Pixel-wise randomization only entails initializing guardian map of the same size as the input frame and adding it pixel-wise. The experiments are conducted on two widely used medium-sized datasets KoNViD-1k and LIVE-VQC, and the results are given in Tables V and VI. From the above two Tables, the following observations can be made.
@@
1) Pixel-wise randomization exhibits incredible contribution to robustness across all models, yet has minimal impact on model performance. Specifically, compared with the original models, the pixel-wise randomization improves the robustness by a remarkable 107.3% on SRCC averaged on the above two datasets when the performance degradation is only 0.8% on SRCC, which proves the effectiveness and versatility of our proposed strategy. 2) Temporal information extraction also plays a key role in model robustness, and it is the sole strategy that contributes to improving model performance. Specifically, incorporating inter-frame information enhances performance by 4.0% and boosts robustness by 41.4% on SRCC averaged on the above two datasets. This substantiates the significance of inter-frame information in the VQA task from an alternative perspective. 3) Spatial grid sampling also demonstrates remarkable defense effects while substantially improving training efficiency. However, a substantial proportion of the video content is discarded during the sampling process, which leads to notable performance degradation when applied to other models. Specifically, applying spatial sampling to other models brings 21.7% performance degradation on SRCC averaged on the above two datasets.
@@
Besides, it can be seen from Tables V and VI that certain strategies may have a detrimental effect on performance due to the inherent tradeoff between security on adversarial videos and accuracy on clean videos. To ensure the security of SecureVQA as well as its accuracy, we adopt three strategies to optimize this trade-off. 1) We employ spatial grid sampling in the intra-branch to enhance the model robustness. However, this may result in the loss of some global information. Therefore, we leverage the entire frame information in the inter-branch to improve performance on clean videos. 2) We set the value of guardian map to +1 or -1, which can not only effectively mitigate the attack effect on adversarial videos but also have minimal alterations to clean videos. 3) We extract inter-frame information to improve model performance on clean videos, while resisting the impact of intra-frame perturbations on robustness. In summary, the experiments in this part demonstrate the versatility of the proposed framework, showcasing its ease of application to other models. Additionally, it empowers future model designers to select different strategies based on their specific requirements.
@@
Secondly, the mechanism of the pixel-wise randomization through guardian maps to the robustness of the model is further explored. The experiments are conducted via attacking VSFA [17] on the KoNViD-1k. We adopt the only black-box attack method [4], which tentatively perturbs a portion of the frame while leaving the remainder untouched. Therefore, the proportion of the disturbed region (i.e. pink area in Fig. 4) indicates the effectiveness of the pixel-wise randomization for the model’s robustness, i.e. smaller disturbed area corresponds to higher robustness. During the attack process, guardian maps (GMs) are respectively added to the attacked region, non-attack region, and the entire frame to investigate the areas where guardian maps provide effective defense. From Fig. 4, three results can be obtained. 1) From (a) and (b), it can be seen that the guardian map neutralizes the carefully designed adversarial perturbations so that attack is compromised to some extent. 2) From (a) and (c), it can be observed that the most contribution of the guardian map comes from its modifications in non-attack region. This can be explained by the fact that adding guardian map to non-attack region can confuse the attacker whether the perturbations or the guardian map are contributing to the adversarial effect at each query. Besides, it can also provide global immunity to the model, ensuring robustness against local perturbations in advance. As a consequence, the model exposes less weakness to malicious queries from adversarial attackers. 3) From (a) to (d), the phenomenon described in the above two cases coexists when guardian map is added to the entire frame. It further diminishes the disturbed area, and thus disables the attack effect, attaining a desirable result close to the score before the attack.
@@
Thirdly, some methods like FAST-VQA [18] also adopt spatial grid sampling. In Table VII, we verify its defense effect compared with SecureVQA. It can be observed that SecureVQA exhibits greater robustness compared with [18]. Specifically, SecureVQA improves the average robustness by 4.7% and 3.6% on SRCC and PLCC compared with [18]. This can be attributed to the defensive capabilities of the strategies like guardian map and inter-frame information in SecureVQA. Furthermore, the primary contribution of this paper is not spatial grid sampling, as only adopting guardian map can provide better defense effects (as shown in Tables V and VI). The primary concern of this paper is to design and discover some strategies tailored to the VQA and can be easily integrated into existing models, enabling the model to exhibit defense capabilities without compromising performance on clean videos. Additionally, different strategies in SecureVQA have different characteristics (as shown in Tables V and VI), which empowers future model designers to select different strategies based on their specific requirements.

### v2

@@ Abstract
The exponential surge in video traffic has intensified the imperative for Video Quality Assessment (VQA). Leveraging cutting-edge architectures, current VQA models have achieved human-comparable accuracy. However, recent studies have revealed the vulnerability of existing VQA models against adversarial attacks. To establish a reliable and practical assessment system, a secure VQA model capable of resisting such malicious attacks is urgently demanded. Unfortunately, no attempt has been made to explore this issue. This paper first attempts to investigate general adversarial defense principles, aiming to endow existing VQA models with security. Specifically, we first introduce random spatial grid sampling on the video frame for intra-frame defense. Then, we design pixel-wise randomization through a guardian map, globally neutralizing adversarial perturbations. Meanwhile, we extract temporal information from the video sequence as compensation for inter-frame defense. Building upon these principles, we present a novel VQA framework from a security-oriented perspective, termed SecureVQA. Extensive experiments indicate that SecureVQA sets a new benchmark in security while achieving competitive VQA performance compared with state-of-the-art models. Ablation studies delve deeper into analyzing the principles of SecureVQA, demonstrating their generalization and contributions to the security of leading VQA models. The code is available at https://github.com/GZHU-DVL/SecureVQA.
@@ I Introduction
The tremendous surge of streaming media platforms like YouTube, Facebook, and TikTok has resulted in an exponential increase in video traffic [1]. This phenomenon has transformed ordinary individuals into both providers and end-users of video services. There is an essential need for Video Quality Assessment (VQA) to automatically evaluate the quality of videos and detect videos that are not up to standard [2, 3]. In general, VQA is categorized into three types based on the availability of the pristine video: Full-Reference VQA (FR-VQA), Reduced-Reference VQA (RR-VQA), and No-Reference VQA (NR-VQA). Due to the characteristics of videos on the Internet, NR-VQA has garnered the widest attention [4, 5, 40, 41, 44]. In recent years, driven by the development of cutting-edge architectures, NR-VQA has experienced rapid advancements and achieved human-comparable performance [45, 46, 22].
@@
With the above considerations, this paper makes the first attempt to investigate some general adversarial defense principles that can be applied to existing VQA models in a plug-and-play manner and design a robust NR-VQA framework from a security standpoint, termed SecureVQA. It is worth noting that SecureVQA does not simply adopt the defense strategies used in classification. Instead, it ingeniously devises solutions tailored to the characteristics of the VQA task. In summary, the key idea behind the proposed framework is to weaken the adversarial effect by subtly introducing some randomness and part of additional information as a guarantee. Specifically, SecureVQA designs two branches and applies three defense strategies to each branch for this purpose, each of which can serve as a general component for existing VQA models.
Firstly, we empirically deem randomness as the natural enemy against optimization-based adversarial attacks. Based on it, we introduce random spatial grid sampling on the video frame for intra-frame defense, which can improve the model’s robustness while preserving the detailed information of the video frame. Secondly, a universal strategy termed pixel-wise randomization is proposed. The pixel-wise randomization deliberately introduces random imperceptible modifications through a guardian map to the video before assessment, effectively destroying the carefully crafted perturbations while maximally avoiding the loss of intra-frame information of the video. Overall, the intra-frame defense aims to disrupt adversarial perturbations within individual frames, thereby improving the model’s security. Thirdly, incorporating additional temporal information is also crucial for model robustness. Given that existing attack methods primarily focus on introducing spatial perturbations, the distinctive temporal information of videos remains unaffected in the intra-frame attacks. Therefore, we introduce inter-frame defense to resist intra-frame attacks. The objective of SecureVQA against adversarial attacks is shown in Fig. 1. It is anticipated that our framework can not only have strong anti-interference ability for carefully designed adversarial videos but also obtain superior VQA performance. The major contributions of this work are summarized as follows:
@@ II-A NR-VQA Models
Since [21] first introduced deep learning into NR-VQA, significant progress has been made in improving the capabilities of models from multiple perspectives [38, 39, 45, 46, 22]. [23] proposed an NR-VQA method based on spatiotemporal feature fusion and hierarchical information integration through deep neural networks. [52] considered the influence of motion distortion and processed the video into continuous chunks, enabling the extraction of spatial-temporal features from the video. [28] introduced texture information as a complementary component to content features, aiming to compensate for the human eye’s tendency to overlook distortions that occur in regions with intricate textures. [25] incorporated a blur encoder into the model to detect the blur effect, recognizing the significant influence of video shakiness on quality. [27] discovered that abrupt scene transitions result in additional degradation in quality, and the Human Visual System (HVS) tends to allocate different levels of attention to frames with varying contents. Subsequently, two Transformer-based modules were designed to handle these two characteristics. [29] systematically divided the video into 13 dimensions to perform a detailed analysis of the individual factors influencing video quality. Additionally, a vision-language-based model was designed to jointly predict specific quality factors and overall quality scores. [24] conducted a thorough survey of HVS, and designed five modules to mimic the five key characteristics of HVS. [26] partitioned the VQA model into technical and aesthetic branches, allowing for a comprehensive evaluation of video quality from two distinct perspectives. [15] explored the potential of Large Language Models (LLMs) in VQA tasks for the first time. However, the results indicated that it was extremely difficult for LLMs to output accurate scores directly. Subsequently, [16] transformed score regression into a classification problem and used the weighted sum as the output scores, validating the ability of LLMs in VQA.
@@ II-B Adversarial Attacks on QA models
In the past two years, some studies have revealed that, similar to classification models, QA models are also vulnerable to adversarial attacks [6, 7, 8, 14, 19, 20]. [6] made the first attempt to attack the NR-IQA models under the white-box setting. The perturbation was randomly initialized from a discrete set. During the optimization process, the steepest ascent method was utilized to calculate the gradient and update the perturbation at each iteration. To ensure the visual quality of the adversarial images, [6] performed human subjective experiments to select the adversarial image that is most similar to the original one. [7] aimed to increase the quality scores of all QA models by training a fixed-size universal perturbation. In their loss function, a normalization factor was introduced to balance the trade-off between the attack effect and the visual quality. [8] conducted the first endeavor to attack the NR-VQA models under the black-box setting, where the attack problem was formulated as misleading the adversarial video’s estimated quality score far away from its ground-truth score towards a specific boundary. Recognizing the time-consuming and labor-intensive human subjective experiments, they employed objective metrics $L_{2}$ and $L_{\infty}$ norm to constrain the magnitude of perturbation automatically. Besides, [13] proposed a black-box adversarial attack against NR-IQA models through a random search method in a greedy way. [17] proposed the correlation-error-based attack framework, which specifically perturbed the score variation on individual images and the correlation within the image set. [11] proposed a query-based black-box attack method against NR-IQA models with adaptive iterative attacks guided by the initial attack direction.
@@ III-A Intra-frame defense
which can effectively eliminate the carefully crafted adversarial perturbation while preserving the visual quality of the defense video with minimal alterations. After applying the spatial sampling and pixel-wise randomization strategies, the input video $\mathbf{V}$ is transformed into $\mathbf{F}^{\text{intra}}$ in the intra-frame branch, and then $\mathbf{F}^{\text{intra}}$ is fed into video Swin-Transformer (Swin-T) to obtain intra-embeddings $\mathbf{E}^{\text{intra}}$ as
@@ III-B Inter-frame defense
where $s:s+d$ denotes selecting $d$ consecutive frames for extracting temporal information, and $s$ is set to a different value from the intra-frame branch. Considering that the details and high-frequency noise of the video frame are preserved through spatial grid sampling in the intra-frame branch, we therefore employ the resize operation as a supplementary in the inter-frame branch to comprehensively extract video embeddings by
@@
Similarly, the pixel-wise randomization is also adopted through the guardian map to introduce randomness and neutralize the meticulously crafted perturbation due to its minimal impact on performance and significant contribution to security by
@@
Temporal information extraction: After applying the resize operation and pixel-wise randomization strategies, the input video $\mathbf{V}$ is transformed into $\mathbf{F}^{\text{inter}}$ in the inter-frame branch, and then $\mathbf{F}^{\text{inter}}$ is fed into SlowFast by
@@
TABLE I: Performance evaluations under white-box setting on VSFA [21], MDTVSFA [44], TiVQA [28], BVQA-2022 [42], and HVS-5M [24]. Here, the performance before the attack is marked in gray.
@@
TABLE II: Performance evaluations under black-box setting on VSFA [21], MDTVSFA [44], TiVQA [28], BVQA-2022 [42], and HVS-5M [24]. Here, the performance before the attack is marked in gray.
@@ IV-A Experimental setup
Attack strategy: Currently, research on the security of VQA is in its preliminary stage. Most of the related work focus on adversarial attacks on IQA. As far as we know, [8] is the only representative adversarial attack on VQA under the black box setting. Therefore, we select [8] as the white-box and black-box attack benchmark in this paper, which designs an efficient way to address the computation issue within adversarial video. Specifically, [8] employs the Projected Gradient Descent (PGD) [53] algorithm for white-box attack, restricting the pixel-wise $L_{2}$ or $L_{\infty}$ norm of the perturbations to 1/255 or 3/255 in each iteration. In a black-box attack, [8] tentatively perturbs a portion of the frame (56$\times$56 pixels) while leaving the remainder untouched at each query. Such perturbations would be kept if the model’s score changes in the expected directions, and abandoned otherwise. After 300 times of queries, the resulting adversarial frame is successfully obtained. Besides, considering the computational complexity, [8] selects 50 videos in the dataset to verify the model security. In this paper, we adopt the same setting as [8] for fair comparison.
Implementation details: The experiments are built on Pytorch 3.8 and run on one RTX 4090 GPU card. In the two branches that handle the intra-frame and inter-frame information, the Swin-Transformer [30] pre-trained on ImageNet [43] and SlowFast [35] pre-trained on Kinetics [36] are utilized as backbone networks to extract embeddings. Wherein, the parameters of Swin-Transformer are set to be trainable, and the parameters of SlowFast are set to be frozen due to its large amount of computation. Such a training setup can speed up the training process. In the skip sampling and continuous sampling process, $n$ and $d$ (in Equations 1 and 7) are set to 2 and 32 respectively, and $\mathbf{V}_{t}^{\text{inter}}$ is resized to 224 $\times$ 224 (in Equation 8). In addition, the inter-frame branch splits the video into 16 segments. In the spatial sampling process, $G$ and $S$ (in Equations 2 and 3) are set to 7 and 32, respectively, as done in [22]. The Adam optimizer with the initial learning rate of 0.001 and batch size 12 is used for training.
TABLE III: Evaluations on the large-scale dataset. Here, the top three performers are highlighted in boldface, underlined, and wavy line, respectively, in descending order of performance.
@@
TABLE IV: Evaluations on smaller datasets. Here, the top three performers are highlighted in boldface, underlined, and wavy line, respectively, in descending order of performance.
@@ IV-B Evaluations on model security
To further validate the security of SecureVQA and its ability to maintain the imperceptibility of the perturbations in the original video, we present some examples of the predicted scores of the original video frames and the corresponding adversarial video frames in Fig. 3. It can be observed that SecureVQA offers formidable security against both white-box and black-box attacks. Specifically, SecureVQA not only prevents high-quality videos from being rated as low quality, it also works the other way around. Furthermore, the modifications introduced by SecureVQA are virtually undetectable, preserving the detailed information and the quality of the video content to the fullest extent.
TABLE V: Ablation studies on KoNViD-1k. Here, the performance before the attack is marked in gray, and the performance improvement and degradation are colored in red and blue, respectively.
@@
TABLE VI: Ablation studies on LIVE-VQC. Here, the performance before the attack is marked in gray, and the performance improvement and degradation are colored in red and blue, respectively.
@@ IV-C Evaluations on quality prediction
In this part, the performance of SecureVQA on clean videos is evaluated on the large-scale LSVQ dataset. Besides, it is further fine-tuned on the small-scale datasets, and the results are given in Tables III and IV, respectively. It can be seen that SecureVQA achieves competitive performance compared with SOTA VQA models. Specifically, SecureVQA has consistently secured a position within the top three across all test categories. Meanwhile, its remarkable performance on the LIVE-VQC dataset, which encompasses rich motion information, can be attributed to SecureVQA’s meticulous consideration of inter-frame information and its effective fusion of embeddings from two distinct branches. Considering that SecureVQA takes security as a prerequisite and then ensures its exceptional VQA performance. As a result, certain strategies like the guardian map may have a detrimental effect on performance due to the inherent tradeoff between security and accuracy (as shown in the ablation study). Despite this, SecureVQA exhibits an accuracy difference of merely 1% to 2% compared to the most advanced DOVER [26]. The results from both this part and the previous part confirm that SecureVQA has successfully achieved the expected goal depicted in Fig. 1. It not only demonstrates a formidable resistance against adversarial attacks but also attains superior VQA performance.
@@ IV-D Ablation study
Figure 4: The defense effect of the guardian map added in different regions and corresponding predicted scores. Here, the pink area denotes the size of the disturbed area after the attack.
Firstly, the generalization of these three strategies to existing models is verified. Each strategy is individually incorporated into the original model in a plug-and-play manner to verify its defense effect and characteristics. Specifically, 1) Inter-frame embeddings can be easily concatenated with the embeddings of existing models. 2) The existing models typically feed pre-processed frames into the networks, thus we simply need to substitute them to the frames processed by spatial grid sampling. 3) Pixel-wise randomization only entails initializing the guardian map of the same size as the input frame and adding it pixel-wise. The experiments are conducted on two widely used medium-sized datasets KoNViD-1k and LIVE-VQC, and the results are given in Tables V and VI. From the above two Tables, the following observations can be made.
@@
1) Pixel-wise randomization exhibits incredible contribution to robustness across all models, yet has minimal impact on model performance. Specifically, compared with the original models, the pixel-wise randomization improves the robustness by a remarkable 107.3% on SRCC averaged on the above two datasets, while the performance degradation is only 0.8% on SRCC, demonstrating the effectiveness and versatility of our proposed strategy. 2) Temporal information extraction also plays a key role in model robustness, and it is the sole strategy that contributes to improving model performance. Specifically, incorporating inter-frame information enhances performance by 4.0% and boosts robustness by 41.4% on SRCC averaged on the above two datasets. This substantiates the significance of inter-frame information in the VQA task from an alternative perspective. 3) Spatial grid sampling also demonstrates remarkable defense effects while substantially improving training efficiency. However, a substantial proportion of the video content is discarded during the sampling process, which leads to notable performance degradation when applied to other models. Specifically, applying spatial sampling to other models brings 21.7% performance degradation on SRCC averaged on the above two datasets.
@@
Besides, it can be seen from Tables V and VI that certain strategies may have a detrimental effect on performance due to the inherent tradeoff between security on adversarial videos and accuracy on clean videos. To ensure the security of SecureVQA as well as its accuracy, we adopt three strategies to optimize this trade-off. 1) We employ spatial grid sampling in the intra-branch to enhance the model’s robustness. However, this may result in the loss of some global information. Therefore, we leverage the entire frame information in the inter-branch to improve performance on clean videos. 2) We set the value of guardian map to +1 or -1, which can not only effectively mitigate the attack effect on adversarial videos but also have minimal alterations to clean videos. 3) We extract inter-frame information to improve model performance on clean videos, while resisting the impact of intra-frame perturbations on robustness. In summary, the experiments in this part demonstrate the versatility of the proposed framework, showcasing its ease of application to other models. Additionally, it empowers future model designers to select different strategies based on their specific requirements.
@@
Secondly, the mechanism of the pixel-wise randomization through guardian maps to the robustness of the model is further explored. The experiments are conducted via attacking VSFA [21] on the KoNViD-1k. We adopt the only black-box attack method [8], which tentatively perturbs a portion of the frame while leaving the remainder untouched. Therefore, the proportion of the disturbed region (i.e. pink area in Fig. 4) indicates the effectiveness of the pixel-wise randomization for the model’s robustness, i.e. smaller disturbed area corresponds to higher robustness. During the attack process, guardian maps (GMs) are respectively added to the attacked region, the non-attack region, and the entire frame to investigate the areas where guardian maps provide effective defense. From Fig. 4, three results can be obtained. 1) From (a) and (b), it can be seen that the guardian map neutralizes the carefully designed adversarial perturbations so that the attack is compromised to some extent. 2) From (a) and (c), it can be observed that the most significant contribution of the guardian map comes from its modifications in the non-attack region. This can be explained by the fact that adding a guardian map to a non-attack region can confuse the attacker whether the perturbations or the guardian map are contributing to the adversarial effect at each query. Besides, it can also provide global immunity to the model, ensuring robustness against local perturbations in advance. As a consequence, the model exposes fewer weaknesses to malicious queries from adversarial attackers. 3) From (a) to (d), the phenomenon described in the above two cases coexists when the guardian map is added to the entire frame. It further diminishes the disturbed area, and thus disables the attack effect, attaining a desirable result close to the score before the attack.
@@
Thirdly, some methods like FAST-VQA [22] also adopt spatial grid sampling. In Table VII, we verify its defense effect compared with SecureVQA. It can be observed that SecureVQA exhibits greater robustness compared with [22]. Specifically, SecureVQA improves the average robustness by 4.7% and 3.6% on SRCC and PLCC compared with [22]. This can be attributed to the defensive capabilities of strategies like the guardian map and inter-frame information in SecureVQA. Furthermore, the primary contribution of this paper is not spatial grid sampling, as only adopting the guardian map can provide better defense effects (as shown in Tables V and VI). The primary concern of this paper is to design and discover some strategies tailored to the VQA and can be easily integrated into existing models, enabling the model to exhibit defense capabilities without compromising performance on clean videos. Additionally, different strategies in SecureVQA have different characteristics (as shown in Tables V and VI), which empower future model designers to select different strategies based on their specific requirements.